OpenStudy (anonymous):

how do u solve 14b=16(b+12) show work step by step

OpenStudy (anonymous):

You should distribute 16 into (b+12). after that get all the b's on one side and then divide to get b on its own.

OpenStudy (anonymous):

14b=16(b+12) First, use the distributive property 14b=16b+192 Now, subtract by 16b -2b = 192 Divide by -2 b=-96

OpenStudy (merchandize):

14b=16b+192 14b-16b=192 -2b=192 b=-192/2 b=-96